### v1

# Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching

Abstract
Entity matching (EM) is a critical step in entity resolution. Recently, entity matching based on large language models (LLMs) has shown great promise. However, current LLM-based entity matching approaches typically follow a binary matching paradigm that ignores the global consistency between different records. In this paper, we investigate various methodologies for LLM-based entity matching that incorporate record interactions from different perspectives. Specifically, we comprehensively compare three representative strategies: matching, comparing, and selecting, and analyze their respective advantages and challenges in diverse scenarios. Based on our findings, we further design a compositional entity matching (ComEM) framework that leverages the composition of multiple strategies and LLMs. In this way, ComEM can benefit from the advantages of different sides and achieve improvements in both effectiveness and efficiency. Experimental results show that ComEM not only achieves significant performance gains on various datasets but also reduces the cost of LLM-based entity matching in real-world scenarios.
[…]
Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching
Tianshu Wang1,2, Hongyu Lin1, Xiaoyang Chen3, Xianpei Han1,4, Hao Wang5, Zhenyu Zeng5, Le Sun1,4 1 Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2 Hangzhou Institute for Advanced Study, University of Chinese Academy of Sciences 3 University of Chinese Academy of Sciences 4 State Key Laboratory of Computer Science, Institute of Software, Chinese Academy of Sciences 5 Alibaba Cloud Group {tianshu2020, hongyu, xianpei, sunle}@iscas.ac.cn, chenxiaoyang19@mails.ucas.ac.cn, cashenry@126.com, zhenyu.zzy@alibaba-inc.com
1 Introduction
Figure 1: Three strategies for EM with LLMs.
Entity resolution (ER), also known as record linkage (Fellegi and Sunter, 1969) or deduplication (Elmagarmid et al., 2007), aims to identify and merge records that refer to the same real-world entity. Entity matching (EM) is a critical step in entity resolution that uses complex techniques to identify matching records from candidate pairs filtered by the blocking step (Papadakis et al., 2021). The recent emergence of large language models (LLMs) has introduced a new zero- or few-shot paradigm to EM, showing great promise (Narayan et al., 2022; Peeters and Bizer, 2023b, a; Fan et al., 2023; Li et al., 2024). LLM-based entity matching methods can achieve similar or even better performance than deep learning methods trained on large amounts of data, and are less susceptible to the unseen entity problem (Wang et al., 2022; Peeters et al., 2024).
However, current LLM-based entity matching methods identify matches by classifying each pair of records independently. This matching strategy ignores the global consistency between records and thus leads to suboptimal results. On the one hand, ER applications require more than independent classification due to the interconnected nature of record relationships (Getoor and Machanavajjhala, 2012). For example, in record linkage (a.k.a. clean-clean ER), a single record from one data source typically matches at most one record from another data source, because there are usually no duplicates in a single database. Unfortunately, matching-based approaches do not take advantage of this nature of record linkage well. On the other hand, this strategy ignores the capabilities of LLMs to handle multiple records simultaneously to distinguish similar records. Using the records in Figure 1(c) as an example, if “Cruzer Glide”, “Cruzer Force”, and “Extreme Pro” appear in different records of the same context, LLMs are more likely to recognize that these are different models of SanDisk flash drives and use them as the basis for identifying matches. As a result, the matching strategy cannot fully unleash the potential of LLMs in EM.
In this paper, we thoroughly investigate three strategies for entity matching with LLMs that incorporate record interactions from different perspectives, as shown in Figure 1. Specifically, apart from the conventional matching strategy shown in Figure 1(a), we investigate two additional strategies that leverage information from other records: 1) the comparing strategy, which identifies the record out of two that is more likely to match the anchor record, as shown in Figure 1(b); 2) the selecting strategy, which directly chooses the record from a list that is most likely to match the anchor record, as shown in Figure 1(c). Our research suggests that for entity matching with LLMs, incorporating record interactions is critical and can significantly improve EM performance in various scenarios. Therefore, the global selecting strategy is often the most effective. However, we also observe that the selecting strategy suffers from the position bias and limited long context understanding of current LLMs, which restricts its generality in applications.
Based on our findings, we design a compositional entity matching (ComEM) framework that leverages the composition of multiple strategies and LLMs for both effectiveness and efficiency. Specifically, given an entity record and its $n$ candidate matches, we first preliminarily rank and filter these candidates using the local matching or comparing strategy, implemented with a medium-sized LLM. We then perform fine-grained identification on the top $k$ candidates using the global selecting strategy, facilitated by a powerful LLM. This approach not only mitigates the challenges and biases faced by the selecting strategy with too many options, but also reduces the cost of LLM invocations caused by composing multiple strategies. Consequently, by integrating the advantages of different strategies and LLMs, ComEM achieves a more effective and efficient entity matching process.
To investigate different strategies with LLMs and to evaluate our ComEM framework, we conducted in-depth experiments on 8 ER datasets. Experimental results verify the effectiveness of incorporating record interactions through the selecting strategy, with an average 13.39% improvement in F1 over the current matching strategy. In addition, we examined the effect of 10 different LLMs using these strategies on identification and ranking. Based on the results, our ComEM framework is able to further improve the average F1 of the single selecting strategy by up to 8.46% while reducing the cost.
Contributions.
Generally speaking, our contributions can be summarized as follows:
  • We investigate three strategies for entity matching with LLMs, and delve into their advantages and shortcomings in different scenarios.
  • We design a ComEM framework by integrating the advantages of different strategies and LLMs to address the challenges of EM.
  • We conduct thorough experiments to investigate these strategies for EM and verify the effectiveness of our proposed framework.
2 Related Work
2.1 Entity Resolution
As the core of data integration and cleaning, entity resolution has received extensive attention over the past six decades (Fellegi and Sunter, 1969; Getoor and Machanavajjhala, 2012; Binette and Steorts, 2020; Papadakis et al., 2021). The blocking-and-matching pipeline has become the mainstream of entity resolution, where blocking filters out obviously dissimilar records and matching identifies duplicates through complex techniques.
Blocking.
Traditional blocking approaches group records into blocks by shared signatures, followed by cleaning up unnecessary blocks and comparisons (Papadakis et al., 2022). Meta-blocking further reduces superfluous candidates by weighting candidate record pairs and graph pruning (Papadakis et al., 2014). Recently, nearest-neighbor search techniques, especially cardinality-based, have gained more attention and achieved the state-of-the-art results for blocking (Thirumuruganathan et al., 2021; Paulsen et al., 2023).
Entity Matching.
The open and complex nature of entity matching has spurred the development of a variety of approaches to address this persistent challenge, including rule-based (Benjelloun et al., 2009; Li et al., 2015), distance-based (Bilenko et al., 2003), and probabilistic methods (Fellegi and Sunter, 1969; Wu et al., 2020), etc. With the introduction of deep learning methods (Mudgal et al., 2018), especially pre-trained language models (Li et al., 2020), entity matching has made significant progress (Barlaug and Gulla, 2021; Tu et al., 2023). The emergence of LLMs brings a new zero- or few-shot paradigm to entity matching (Narayan et al., 2022; Peeters and Bizer, 2023a), alleviating training data requirements and the unseen entity problem (Wang et al., 2022; Peeters et al., 2024). Most deep learning and LLM-based approaches treat entity matching as an independent classification problem, except for GNEM (Chen et al., 2021) which models this task as a collective classification task on graphs. To the best of our knowledge, this is the first effort to formulate entity matching as a comparison or selection task with LLMs.
2.2 Large Language Model
The advent of LLMs like ChatGPT marks a significant advance in artificial intelligence, offering unprecedented natural language understanding and generation capabilities, and even general intelligence (Bubeck et al., 2023). By scaling up the model and data size of pre-trained language models, LLMs exhibit emergent abilities (Wei et al., 2022) and can thus solve a variety of complex tasks by “prompt engineering” without “fine-tuning”. For more technical details on LLMs, we refer the reader to the related survey (Zhao et al., 2023).
3 Problem Formulation
In this paper, we formulate the task of entity matching as the process of identifying matching records from a given entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$. This formulation mitigates the limitations of independent pairwise matching and responds to real-world entity resolution scenarios. Current state-of-the-art blocking methods adhere to the k-nearest neighbor (kNN) search paradigm, which retrieves a list of candidate matches for each entity record by transforming it into a sparse or dense vector and performing the kNN search. In addition, this formulation accommodates both single-source deduplication and dual-source record linkage, and makes good use of the 1-1 assumption, i.e., record $r$ matches at most one of the candidate matches $R_{c}$. This assumption is widespread in record linkage, and deduplication with canonical entity construction.
4 Entity Matching with LLMs
In this section, we first introduce three strategies for entity matching with LLMs. We then propose our ComEM framework that leverages the composition of multiple strategies and LLMs.
4.1 LLM as a Matcher
Recent work formulates entity matching as a binary classification task based on LLMs (Narayan et al., 2022; Peeters and Bizer, 2023b, a; Fan et al., 2023; Li et al., 2024). In this strategy, an LLM acts as a pairwise matcher to determine whether two records match. Specifically, given an entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$, this approach independently classifies each pair of records $(r,r_{i})_{1\leq i\leq n}$ as matching or not by interfacing LLMs with an appropriate matching prompt $\operatorname{P}_{m}$, as shown in Figure 1(a):
$\operatorname{LLM}_{m}\colon\left\{(r,r_{i})\mid r_{i}\in R_{c}\right\}\xrightarrow{\operatorname{P}_{m}}\left\{\mathrm{Yes},\mathrm{No}\right\}$
Unlike previous studies, the core of LLM-based applications is to prompt LLMs to generate the correct answer, namely prompt engineering. An appropriate prompt should include the task instruction, such as “Do these two records refer to the same real-world entity? Answer Yes or No.” Optionally, a prompt could include detailed rules or several in-context learning examples to guide LLMs in performing this task. Given the need for long contexts in other strategies, and the instability of existing prompt engineering methods for entity matching Peeters and Bizer (2023a), we only attempt few-shot prompting on the matching strategy and leave the exploration of better prompt engineering with different strategies to future work.
This independent matching strategy ignores the global consistency of ER, as well as the capabilities of LLMs to incorporate record interactions. For record linkage, according to the well-known 1-1 assumption Getoor and Machanavajjhala (2012), each entity record $r$ matches at most one record of the candidate matches $R_{c}$. For deduplication, this restriction of at most one match per record can also be satisfied by constructing canonical entities. The traditional solution to satisfy these constraints is to construct a graph based on the similarity scores $s_{i}$ of record pairs $(r,r_{i})$ and to further cluster on the similarity graph. We can obtain the similarity scores from LLMs by calibrating the generated probabilities $p$ of labels Qin et al. (2023). Formally, the similarity score $s_{i}$ can be defined as:
$s_{i}=\begin{cases}1+p(\mathrm{Yes}\mid\operatorname{P}_{m}(r,r_{i})),&\text{if generate ``Yes''}\\ 1-p(\mathrm{No}\mid\operatorname{P}_{m}(r,r_{i})),&\text{if generate ``No''}\end{cases}$
Unfortunately, the generation probabilities are not available for many black-box commercial LLMs. Moreover, the probabilities on short-form labels are misaligned for common open-source chat-tuned LLMs because they are fine-tuned to respond in detail. The need to investigate better strategies for entity matching with LLMs arises in ER.
4.2 LLM as a Comparator
In this section, we introduce a comparing strategy for LLM-based entity matching that simultaneously compares two candidate matches to a given record. Specifically, given an entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$, the comparing strategy compares two records $r_{i}$ and $r_{j}$ from candidate matches $R_{c}$ to determine which is more consistent with record $r$ by interfacing LLMs with a comparison prompt $\operatorname{P}_{c}$, as shown in Figure 1(b):
$\operatorname{LLM}_{c}\colon\left\{(r,r_{i},r_{j})\mid r_{i,j}\in R_{c}\right\}\xrightarrow{\operatorname{P}_{c}}\left\{\mathrm{A},\mathrm{B}\right\}$
where $\mathrm{A}$ and $\mathrm{B}$ are the labels corresponding to record $r_{i}$ and $r_{j}$. Since LLMs are sensitive to the prompt order, we compare each pair of records $(r_{i},r_{j})$ to record $r$ twice by swapping their order, i.e., $\operatorname{LLM}_{c}(r,r_{i},r_{j})$ and $\operatorname{LLM}_{c}(r,r_{j},r_{i})$.
Compared to the matching strategy, this comparing strategy introduces an additional record for more record interactions and shifts the task paradigm. It focuses on indicating the relative relationship between two candidate matches of a given record, rather than making a direct match or no match decision. Therefore, this strategy is suitable for ranking and fine-grained filtering to determine the most likely records for identification.
To rank candidate records using the comparing strategy, we can compute similarity scores to estimate how closely each candidate matches the anchor record. Unlike the matching strategy, the comparing strategy can obtain similarity scores of record pairs using black-box LLMs, which do not provide probabilities. In such case, the similarity score $s_{i}$ of record pair $(r,r_{i})$ can be defined as:
$s_{i}=2\times\sum_{j\neq i}\mathbb{1}_{r_{i}>_{r}r_{j}}+\sum_{j\neq i}\mathbb{1}_{r_{i}=_{r}r_{j}}$
where $\mathbb{1}_{r_{i}>_{r}r_{j}}$ and $\mathbb{1}_{r_{i}=_{r}r_{j}}$ indicate that record $r_{i}$ wins twice and once in comparison with record $r_{j}$ to record $r$. When LLMs do provide probabilities, the similarity scores $s_{i}$ can be defined as:
$s_{i}=\sum_{j\neq i}\left(p(\mathrm{A}\mid\operatorname{P}_{c}(r,r_{i},r_{j}))+p(\mathrm{B}\mid\operatorname{P}_{c}(r,r_{j},r_{i}))\right)$
Here, probabilities are incorporated to calculate similarity scores, integrating a more nuanced assessment from the comparisons made by LLMs.
However, the advantage of the comparing strategy in obtaining similarity scores comes at the cost of using LLMs as the basic unit of comparison and $\mathcal{O}(n^{2})$ complexity. Fortunately, for entity matching, we only care about a small number of most similar candidates, and there are many comparison sort algorithms available to find top-$k$ elements efficiently. In this paper, we use the bubble sort algorithm to find the top-$k$ elements, optimizing the complexity of the comparing strategy to $\mathcal{O}(kn)$. To avoid confusion, we refer to the comparison of all pairs as comparingall-pair in experiments.
4.3 LLM as a Selector
In this section, we propose a selecting strategy that uses an LLM to select the matching record of a given record from a list of candidate matches. Specifically, given an entity record $r$ and its candidate matches $R_{c}=\{r_{1},r_{2},\ldots,r_{n}\}$, this strategy directly selects the match of record $r$ from $R_{c}$ by interfacing LLMs with an appropriate selection prompt $\operatorname{P}_{s}$, as shown in Figure 1(c):
$\operatorname{LLM}_{s}\colon\left\{(r,R_{c})\right\}\xrightarrow{\operatorname{P}_{s}}\left\{1,2,\ldots,n\right\}$
where $1,\ldots,n$ indicates the corresponding record.
In this way, LLMs can be explicitly required to identify only one match per record $r$ in the candidate matches $R_{c}$. Furthermore, feeding LLMs all candidate matches in the same context at a time allows LLMs to make better decisions by considering interactions between candidate records. For example, if “SanDisk Cruzer Glide”, “SanDisk Cruzer Force”, and “SanDisk Extreme Pro” appear in different records of the same context, it is easier for LLMs to recognize that they are different model names of SanDisk flash drives and make the correct choice for the actual match.
One challenge in applying the selecting strategy to LLM-based entity matching is that there is often no actual match of record $r$ in candidate matches $R_{c}$, because entity matching is an extremely imbalanced task. A trivial solution to this challenge could be to perform a pairwise matching after the selection, which would undermine the advantages of the selecting strategy. Another method could be to add “none of the above” as an additional option to allow LLMs to refuse to select any record from the candidate matches. Suppose the selection prompt with the “none of the above” option is $\operatorname{P}_{s_{N}}$, this method can be formulated as:
$\operatorname{LLM}_{s_{N}}\colon\left\{(r,R_{c})\right\}\xrightarrow{\operatorname{P}_{s_{N}}}\left\{0,1,2,\ldots,n\right\}$
where 00 indicates the “none of the above” option.
However, the selecting strategy relies heavily on the capabilities of LLMs for fine-grained understanding and implicit ranking in long contexts. Our experimental results (§ 5.2) show that the current LLMs suffer from position bias, with the selection accuracy decreasing significantly as the position of the matching record in the candidate list increases. In practice, the recall-oriented blocking step often generates a considerable number of candidate matches for each record, exceeding the context length that LLMs can effectively handle. It is a challenge to mitigate the position bias and the long context requirement for the selecting strategy.
4.4 ComEM Framework
Strategy
Similarity
Score
Interaction
Level
# LLM
Invocations
# Prompt
Tokens
Matching $\mathcal{O}(n)$ $\mathcal{O}(1)$
Comparing $\mathcal{O}(n^{2})$ or $\mathcal{O}(n)$ $\mathcal{O}(1)$
Selecting $\mathcal{O}(1)$ $\mathcal{O}(n)$
Table 1: Comparison of different strategies. “” means that the matching strategy can only calibrate similarity scores if the generation probability is available.
Given the advantages and shortcomings of different strategies, we design a compositional entity matching (ComEM) framework. This ComEM framework addresses various challenges of entity matching by integrating the advantages of different strategies and LLMs. Table 1 shows a comparison of these strategies. The matching and comparing strategies are applicable for local ranking, while the selecting strategy is suitable for fine-grained identification. Therefore, as shown in Figure 2, we first utilize a medium-sized LLM to rank and filter the candidate matches $R_{c}$ of record $r$ with the matching or comparing strategy. We then utilize an LLM to identify the match of record $r$ from only top $k$ candidates with the selecting strategy.
Our ComEM framework integrates the advantages of different strategies through a filtering then identifying pipeline. It first utilizes the local matching or comparing strategy to rank candidate matches for preliminary screening, which can effectively mitigate the position bias and the long context requirement of the selecting strategy. It then utilizes the global selecting strategy to incorporate record interactions for fine-grained optimization, which can effectively mitigate the consistency ignorance of the matching strategy. Thus, ComEM is able to strike a balance between entity matching requirements and current LLM capabilities, achieving significant performance improvements.
By integrating LLMs of different sizes, our ComEM framework can also effectively reduce the cost of LLM invocations for entity matching. It is expensive to use commercial LLMs directly in practice because entity matching is a computationally intensive task. Our ComEM framework delegates a significant part of the computation to medium-sized LLMs. Experimental results reveal that the ranking process can be performed well by using open-source medium-sized (3B~11B) LLMs. Therefore, the proper integration of LLMs not only improves the performance of entity matching but also reduces the cost in practice.
Figure 2: Illustration of our ComEM framework.
[…]
Abt-Buy 44.23 29.32 90.00 0.25 64.19 60.00 69.00 0.64 67.94 54.94 89.00 0.09
Amazon-Google 35.47 25.00 61.00 0.24 45.88 55.71 39.00 0.62 53.79 43.29 71.00 0.09
DBLP-ACM 67.35 51.03 99.00 0.35 77.36 73.21 82.00 0.94 70.04 54.80 97.00 0.15
DBLP-Scholar 66.13 55.41 82.00 0.33 62.65 78.79 52.00 0.84 67.40 53.18 92.00 0.12
IMDB-TMDB 65.82 89.66 52.00 0.27 44.12 83.33 30.00 0.70 85.84 78.99 94.00 0.11
IMDB-TVDB 57.59 60.44 55.00 0.24 52.78 86.36 38.00 0.60 72.73 69.72 76.00 0.08
TMDB-TVDB 69.43 55.76 92.00 0.24 80.19 75.89 85.00 0.61 71.60 58.60 92.00 0.08
Walmart-Amazon 41.84 27.16 91.00 0.39 62.56 64.21 61.00 1.03 65.64 53.46 85.00 0.17
Mean 55.98 49.22 77.75 0.29 61.22 72.19 57.00 0.75 69.37 58.37 87.00 0.11
Table 2: Overall performance and cost of different strategies. We bold the best F1 score.
5 Experiments
In this section, we conduct thorough experiments to investigate three strategies for entity matching with LLMs. First, we perform the analysis of different strategies (§ 5.2). Next, we examine the effect of different LLMs on these strategies (§ 5.3). Finally, we verify our ComEM framework that is proposed based on investigation findings (§ 5.4).
5.1 Experimental Setup
Dataset Domain # D1 # D2 # Attr # Pos
Abt-Buy Product 1076 1076 3 1076
Amazon-Google Software 1354 3039 4 1103
DBLP-ACM Citation 2616 2294 4 2224
DBLP-Scholar Citation 2516 61353 4 2308
IMDB-TMDB Movie 5118 6056 5 1968
IMDB-TVDB Movie 5118 7810 4 1072
TMDB-TVDB Movie 6056 7810 6 1095
Walmart-Amazon Electronics 2554 22074 6 853
Table 3: Statistics of experiment datasets.
Datasets.
We focused on the common record linkage that has many open-access datasets. We used 8 ER datasets collected by pyJedAI Nikoletos et al. (2022) because pairwise EM benchmarks Mudgal et al. (2018) for deep learning-based methods involve extensive negative pair downsampling, which conceals real challenges of EM Papadakis et al. (2023). For each dataset with two record collections D1 and D2, we retrieved 10 candidate matches from D2 for each record in D1 using the SOTA blocking method Sparkly Paulsen et al. (2023). The recall of Sparkly on all datasets ranges from 86.57% to 99.96%, demonstrating its effectiveness in retrieving candidate matches. This allows us to investigate in real-world scenarios. Table 3 summarizes the statistics of these datasets.
Evaluation Metrics.
Consistent with prior studies, we report F1, Precision, and Recall on record pairs as evaluation metrics for entity matching. We also report the cost (\$) of LLM invocations [^1] Inference cost is estimated based on the hourly price of the cloud NVIDIA A40..
Implementation Details.
We used ChatGPT (gpt-3.5-turbo-0613) as the main LLM for strategy analysis. For the comparing strategy, we match the top 1 candidate after ranking. We also examined the effect of 9 open-source LLMs on these strategies, including Llama2-7/13/70B-Chat Touvron et al. (2023), Mistral-7B Jiang et al. (2023), Mixtral-8x7B Jiang et al. (2024), Baichuan2-13B-Chat Yang et al. (2023), Vicuna1.5-13B Zheng et al. (2023), SOLAR-10.7B Kim et al. (2023), and Flan-T5-11B Chung et al. (2022). The generation temperature was set to 0 for reproducibility. We sampled 200 records for each dataset, half of which had matches, and formed 2000 pairs of records with their candidate matches. In ComEM, we used Flan-T5-11B to rank candidates with the matching strategy and kept the top 4 candidates for selection.
5.2 Analysis of Strategies
We first compare the performance, cost, and bias of the strategies, with the following findings.
Figure 3: Comparison of different strategies.
Figure 4: Effect of open-source LLMs on different strategies. Star pentagons represent the results of ChatGPT.
Figure 5: F1 score w.r.t. matched record positions.
Finding 1. Incorporating record interactions is essential for entity matching with LLMs. As shown in Table 2, entity matching performance increases as the level of record interaction increases. The comparing strategy outperforms the independent matching strategy by an average of 5.24% F1 score, and the selecting strategy further improves the F1 score by up to 8.15% over the comparing strategy. The selection strategy is optimal in 6 out of 8 datasets, and suboptimal in the remaining two, fully demonstrating the importance of incorporating record interactions. To further verify that these improvements are brought by the strategy, we also perform 6-shot matching, ensuring that the number of records is consistent with the selecting strategy. As shown in Figure 3, the selecting strategy still has a significant F1 advantage over 6-shot matching. These results highlight the importance of incorporating record interactions and open new avenues for entity matching with LLMs.
Finding 2. The selecting strategy is the most cost-effective for entity matching with LLMs. Monetary cost is also an important factor when interfacing LLMs for EM in practice, as it is computationally intensive. As shown in Table 2, the selecting strategy costs less than half of the matching strategy. Although the complexity of the number of prompt tokens for these two strategies is $\mathcal{O}(n)$, the number of times a record is repeatedly prompted to the LLM, as well as the repetition of task instructions, can significantly affect the actual cost. The comparing strategy, however, considers two candidate matches at a time and interfaces the LLM twice, making its cost more than twice that of the matching strategy. Therefore, the selecting strategy stands out for its effectiveness and efficiency.
Finding 3. Strategies that incorporate multiple records suffer from the position bias of LLMs. As shown in Figure 5, the performance of the comparing and selecting strategies decreases significantly as the position of the matching records moves down in the candidate list. For the comparing strategy optimized with bubbling sort, the matching records cannot be ranked at the top if there is any incorrect comparison. LLMs struggle with comparing records, as we will further see in the next section. The selecting strategy also drops nearly 10% in F1, probably due to the limited long context understanding of the LLM. As a result, the position bias of LLMs restricts the performance and generality of the comparing and selecting strategies.
5.3 Effect of LLMs
We further examine the effect of open-source LLMs on these strategies to identify matches or rank.
Finding 4. There is no single LLM that is uniformly dominant across all strategies. Figure 4 shows the efficacy of open-source LLMs on these strategies, with detailed results in Appendix A. We can see that most LLMs perform well in the selecting strategy, but may perform quite poorly in the matching and comparing strategies. This consistent performance of LLMs in the selecting strategy once again confirms the effectiveness of incorporating global information in this way. For the current matching strategy, however, common chat LLMs, especially the Llama2 series and Vicuna1.5, lack the capability to identify matches and produce a large number of false positives. This suggests that LLMs that supervised fine-tuning only on conversation data may not be directly applicable to specific tasks. The situation is even worse for the comparing strategy. Almost all LLMs, even the powerful Mixtral-8x7B, are unable to find more similar records from two candidate matches. These LLMs either make inconsistent choices between two comparisons that exchange the candidate order, or they always output a specific identifier regardless of the records. Although Flan-T5-11B lags behind ChatGPT by about 10% F1 in the selecting strategy, we find that it performs quite well in the matching and comparing strategies. This instruction-tuned LLM follows instructions better than chat-tuned LLMs and can output only the requested labels instead of long-form responses, making it convenient to utilize label generation probabilities. In conclusion, there is a noticeable variance in the capabilities of different LLMs for a single strategy, and the efficacy of different strategies for a single LLM can also be significantly distinct.
Figure 6: Ranking recall@1 w.r.t. model parameters.
Finding 5. Matching strategy is better for ranking and filtering than comparing strategy. The superiority of Flan-T5 in the matching and comparing strategies leads us to explore the possibility of using it to rank and filter candidate matches for the selecting strategy. As shown in Figure 6, the matching strategy outperforms the comparing strategy under different model parameter sizes, even though the latter performs $\mathcal{O}(n^{2})$ comparisons. Their difference is small on the 3B flan-t5-xl and 11B flan-t5-xxl models, but significant on smaller models. This may be due to the fact that these models are trained on many pairwise tasks, such as natural language inference and question answering, but few triple-wise tasks. Therefore, in terms of effectiveness and efficiency, the matching strategy is more suitable for ranking and filtering candidate matches.
5.4 Compositional Entity Matching
We finally verify the feasibility of our ComEM that integrates multiple strategies and LLMs.
Finding 6. By integrating the advantages of different strategies and LLMs, ComEM can accomplish EM more effectively and cost-efficiently. As shown in Figure 7, compared to the optimal selecting strategy using ChatGPT, ComEM achieves a significant F1 improvement (~8%) while spending less. The filtering and identifying pipeline improves precision considerably without sacrificing high recall of the selecting strategy. These results reveal that integrating multiple strategies can complement single strategies and effectively mitigate the position bias of the selecting strategy in long contexts. However, using a single powerful but costly commercial LLM to complete the entire pipeline obscures the cost efficiency of the selecting strategy. By introducing a medium-sized but good at matching LLM for preliminary filtering, ComEM maintains the performance improvement and even spends less than direct selection. As a result, our ComEM framework underscores the importance of task decomposition and LLM composition, illuminating an effective route for compositional entity matching with LLMs.
Figure 7: Performance and cost of compositional entity matching. “M.” and “S.” denote matching and selecting.
Ablation Study.
We perform an ablation study on the number of candidate records for further identification. As shown in Figure 8, recall increases and precision decreases as the number of retained candidate matches increases. Consistent with Figure 5, four is the sweet spot for the selecting strategy with current ChatGPT, which balances precision and recall to achieve the best F1 score.
Figure 8: Average F1, precision, and recall w.r.t. number of candidate matches ($k$) for further identification. See Appendix B for detailed results on each dataset.
6 Conclusion
In this paper, we investigate three strategies for entity matching with LLMs to bridge the gap between local matching and global consistency of ER. Our research shows that incorporating record interactions is essential for LLM-based entity matching. By examining the effect of broad LLMs on these strategies, we further design a ComEM framework that integrates the advantages of multiple strategies and LLMs. The effectiveness and cost efficiency of ComEM highlight the importance of task decomposition and LLM composition, opening up new avenues for entity matching with LLMs.
Limitations
This study aims to investigate different strategies for entity matching with LLMs. We conducted experiments using 1 commercial LLM and 9 open-source LLMs to provide a broad base for our analyses. Nevertheless, the dynamic update of LLMs and the influence of more powerful LLMs need to be further investigated. Besides, prompt engineering or fine-tuned approaches for different strategies are also a worthwhile direction. We have also demonstrated the effectiveness of the composition framework in entity matching that integrates different strategies and LLMs. In future work, we will further uncover the feasibility of automated composition and extend this paradigm to different stages of entity resolution.
Ethical Considerations
We used publicly available LLMs and datasets to conduct experiments in this paper. Our code implementation and application scenarios do not involve any ethical issues.
Appendix A Detailed Results of Open-Source LLMs on Different Strategies
Table 4, Table 5, Table 6, Table 7, Table 8, Table 9, Table 10, Table 11, and Table 12 show the performance of the experimented LLMs with three strategies on all datasets.
Appendix B Ablation Study on Each Dataset
Figure 9 shows the performance of ComEM with respect to the different number of candidates retained for further selection. Similar to Figure 8, recall increases and precision decreases as $k$ increases. For some datasets such as Walmart-Amazon, DBLP-ACM, and TMDB-TVDB, F1 achieves the highest at $k=1$. For some of the other datasets, F1 changes dramatically as $k$ goes from 2 to 5. How to choose $k$ to balance the LLM capabilities and the actual situation is a direction that could be explored. One possible solution might be to dynamically adjust the number according to the probabilities of comparisons for each record.